tell whether the system has one solution infinitely many solutions or no solution

5=1/2x+3y
10-x=6y

Answer by mananth(16946)

5=1/2x+3y
multiply by 2
10=x+6y
6y=-x+10
/6
y=-x/6 +10/6

10-x=6y
/6
y=-x/6 + 10/6
compare with y= mx+b
slopes are same and y intercepts are same
so the lines coincide with each other.
so they have innumerable solution
-